OpenStudy (amonoconnor):

I'm having an issue simplifying and getting the "right" answer that's in my book on this one, and am not sure what I should do differently.... The question is: (x^3 + y^3) = 1; Find y'' by Implicit Differentiation. Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

How I went about solving this one: \[x^3+y^3 = 1\] \[3x^2+3y^2dy/dx = (0)\] \[3x^2=-3y^2dy/dx\] \[dy/dx=\frac{ 3x^2 }{-3y^2}\] \[-x^2/y^2\]

OpenStudy (amonoconnor):

Then, I did the following to find y'': \[y' =-x^2/y^2\] \[= -x^2y^{-2}\] \[y'' = (-2x)y^{-2}+(-x^2)(-2ydy/dx)\]

ganeshie8 (ganeshie8):

Looks good! maybe plugin the value of \(dy/dx\) and simplify

OpenStudy (anonymous):

don't be scared of the quotient rule here either might make it nicer to see what you really have

OpenStudy (amonoconnor):

*Substituting in "(-x^2y^-2)" for dy/dx... \[= -2xy^{-2}+(2x^2y)(-x^2y^{-2})\] \[= -2xy^{-2}+(-2x^4y^{-1})\] \[= -2xy^{-1}(y^3-x^3)\]

OpenStudy (anonymous):

\[-\frac{y^2-2yx^2y'}{y^4}\]

OpenStudy (anonymous):

oops that was wrong

OpenStudy (amonoconnor):

And that's all the farther I know how to go, but my book has: \[\frac{-2x}{y^5}\]

OpenStudy (anonymous):

\[-\frac{2xy^2-2yx^2y'}{y^4}\]

OpenStudy (anonymous):

where you see \(y'\) put \(-\frac{x^2}{y^2}\) then simplify the compound fraction

OpenStudy (anonymous):

\[-\frac{2xy^2-2yx^2y'}{y^4}\\ -\frac{2xy^2+2yx^2\frac{x^2}{y^2}}{y^4}\]

OpenStudy (amonoconnor):

Are you saying that that equals -2x/y^5 ?

OpenStudy (amonoconnor):

I worked through it using the Quotient Rule, but where the heck does the exponent of 5 in the denominator come from?